Write the equation of a line with the slope of -2 which contains the point (3,6) ?? How would i do this?

Answer by mananth(16946)

the general equation for a straight line in the slope intercept form
y=mx+b
m = slope
b= intercept
point (3,6)
slope = -2
plug the values in y= mx +b to find the intersept
6=-2*3+b
6=-6+b
add 6 to both sides
6+6=-6+6+b
12=b
the equation of line with slope -2 and b =6
will be y=-2x+12